OpenStudy (anonymous):

Solve the following for x. (x-2)(3x-5)=0. With explanation please

OpenStudy (hoblos):

x-2=0 => x=2 or 3x-5=0 => x= 5/3

OpenStudy (anonymous):

^^^^^^^^^^^^he is right